# Melding the Serverless Control Plane with the Conventional Cluster Manager for Speed and Resource Efficiency

### v3

@@ Abstract.
Serverless platforms face a trade-off: conventional cluster managers like Kubernetes offer compatibility for co-locating Function-as-a-Service (FaaS) and Backend-as-a-Service (BaaS) components of serverless applications at the cost of high cold-start latency, while specialized FaaS-only systems like Dirigent achieve low latency by sacrificing compatibility, which prevents integrated management and optimization. Our analysis reveals FaaS traffic is bimodal: predictable, sustainable traffic consumes ¿98% of cluster resources, while sporadic excessive bursts stress the control plane’s scaling latency, not its throughput.
@@
With these insights, we design PulseNet, a serverless architecture that employs a dual-track control plane tailored to both traffic types. PulseNet’s standard track manages sustainable traffic with long-lived, full-featured Regular Instances under a conventional cluster manager, preserving compatibility and robust features for the majority of the workload. To handle excessive traffic, an expedited track bypasses the slow manager to rapidly create short-lived, disposable Emergency Instances, minimizing cold-start latency and resource waste from idle instances. This hybrid approach achieves 35% better performance than Dirigent, a FaaS-only system, at the same cost and outperforms other Kubernetes-compatible systems by 1.5–3.5 at a 3–70% lower cost.
1. Introduction
@@
Serverless cloud applications today comprise stateless Function-as-a-Service (FaaS) and conventional stateful Backend-as-a-Service (BaaS) components, organized as DAG workflows. Developers deploy their business logic as functions while offloading the management of the underlying cloud infrastructure – for both FaaS and BaaS – entirely to the providers. The control plane is key to providing high-quality service and resource efficiency when running these applications at the massive cloud scale, which can be achieved by careful placement and load balancing, and timely scaling.
@@
Unfortunately, despite the tight coupling of FaaS and BaaS components in cloud applications, their control planes have no coordination, often rendering co-design and optimizations impractical. For example, AWS Lambda (Agache et al., 2020) and Dirigent (Cvetkovic et al., 2024) manage their own fleets of bare-metal nodes allocated exclusively for FaaS. Other works focus on designs for specific BaaS services (Klimovic et al., 2018; Mahgoub et al., 2021) used by serverless applications, such as storage and database. In contrast, a decade ago, the success of Google’s Borg cluster manager demonstrated the benefits of running various services under the same control plane, achieving better cluster utilization and reducing the operational costs without compromising performance (Verma et al., 2015; Tirmazi et al., 2020). Since then, practitioners have designed many open-source production systems (ope, 2025a; kna, 2025a; fis, 2025; nuc, 2025) running atop Borg’s successor, Kubernetes (k8s, 2025a), which features numerous optimizations and highly-optimized technologies (e.g., intelligent placement, load balancing, networking, and authentication) and can colocate any services in the same cluster. However, state-of-the-art prior works (Cvetkovic et al., 2024; Roy et al., 2022; Singhvi et al., 2021) have concluded that a heavyweight control plane, as in Kubernetes, is too slow to keep up with highly dynamic FaaS workloads (Shahrad et al., 2020), calling for a full revamp of technologies already available for BaaS services running under a conventional control plane.
@@
In contrast, we show that conventional cluster manager’s control plane, although heavyweight, can be seamlessly extended to offer sufficient scaling speed for FaaS workloads, unlocking efficient co-location of FaaS and BaaS systems without compromising performance. Using production Azure Functions traces (Shahrad et al., 2020), we present a thorough analysis of the FaaS workload as well as the performance characteristics in systems, representative of leading commercial offerings (gcr, 2025; aws, 2025c) and open-source systems (Apache, 2025; kna, 2025c). First, we identify two kinds of invocation traffic, sustainable and excessive, showing that the former utilizes 98% of the cluster resources while the latter stresses the control plane but consumes 2% of the cluster resources. Second, we demonstrate that control planes suffer from long instance-creation delays due to frequent interaction with the centralized cluster manager, e.g., when allocating cluster resources and IP addresses, and setting up network routes. However, we show that these systems’ control planes can be tuned to deliver throughput sufficient for a dynamic large-scale FaaS deployment, in contrast to the prior work (Cvetkovic et al., 2024) that has optimized for control-plane throughput. Finally, we identify that current systems waste 70-87% of memory and 9-20% of CPU resources due to prolonged idle instance lifetime and high instance churn, respectively.
@@
Based on the obtained insights, we introduce PulseNet,111We will release the code, traces, and toolchain after publication. a serverless system designed to achieve high performance and low cost while maintaining compatibility with conventional cluster managers, such as Kubernetes (k8s, 2025a). PulseNet employs a novel dual-track control plane comprising standard and expedited tracks. The standard track comprises a Knative-based control plane operating atop a conventional cluster manager that manages long-lived, full-featured Regular Instances for handling predictable, sustainable traffic. This track adjusts the number of instances off the critical path, in cooperation with the underlying cluster manager, hence preserving full compatibility with its rich feature set. Concurrently, the expedited track serves excessive traffic bursts by rapidly creating Emergency Instances that completely bypass the cluster manager and its bookkeeping. The expedited track communicates directly with a node-local agent that spawns single-invocation instances, which the agent shuts down after processing that invocation, minimizing idle instance lifetime and the associated resource waste. These disposable instances are fully compatible with FaaS workloads, albeit they forego the conventional cluster manager’s features for long-running instances, such as readiness and liveliness probes, and advanced networking capabilities – which they do not need – for much faster instance creation.
@@
We prototype PulseNet in vHive (Ustiugov et al., 2021b) as a seamless extension of the Knative control plane running atop unchanged Kubernetes. Evaluation results show that PulseNet provides 35% performance improvement while maintaining cost parity with Dirigent (Cvetkovic et al., 2024). Additionally, when compared with other Kubernetes-compatible alternatives, PulseNet delivers 1.5-3.5 better performance than the systems with a synchronous control plane, as in AWS Lambda (ray, 2025), while reducing the operational cost by 8-70%; and achieves 1.7-3.5 better performance than the systems with an asynchronous control plane, as in Knative, while reducing the cost by 3-65%. Finally, compared with the systems featuring a linear-regression and the state-of-the-art NHITS prediction models (Challu et al., 2023; Joosen et al., 2023), PulseNet delivers up to 4 better performance while reducing costs by 35-40%.
@@
We summarize the comparisons across state-of-the-art systems in Table 1 and illustrate the scaling speed and resource efficiency in Figure 1 on a trace snippet taken from Azure Functions trace (Shahrad et al., 2020). In contrast to the baselines, PulseNet closely follows traffic changes by rapidly creating emergency instances whenever the standard track lags behind and quickly tearing these instances down after the traffic trend settles.
2. Background
We start with an overview of the two approaches for serverless application deployment, namely FaaS and BaaS colocation in the same cluster or disaggregation across different clusters. We then detail the state-of-the-art control plane architectures and their associated trade-offs.
2.1. FaaS and BaaS: Co-located vs. Disaggregated
One deployment option, colocation, implies running both FaaS and BaaS under a unified, feature-rich cluster manager, e.g., Borg (Verma et al., 2015; Tirmazi et al., 2020) or Kubernetes (k8s, 2025a). This approach is popular in many open-source serverless systems (fis, 2025; kna, 2025a; ope, 2025a; Apache, 2025) because it enables integrated management and placement optimizations when colocating various microservices (Zhang et al., 2021; Bazzaz et al., 2025; Verma et al., 2015) and provides support for monitoring (pro, 2025; gra, 2025; ope, 2025b), logging (flu, 2025), network traffic control (cal, 2025), network mounts (k8s, 2025d), and versioned deployments (k8s, 2025b). However, supporting this rich and general-purpose feature set often contributes to long cold-start delays when scaling due to traffic changes in highly dynamic workloads, such as FaaS, whose performance is especially sensitive to cold starts (Shahrad et al., 2020; Cvetkovic et al., 2024).
Alternatively, providers can use separate, specialized cluster managers for FaaS (Cvetkovic et al., 2024; Singhvi et al., 2021; Mvondo et al., 2021; Fuerst and Sharma, 2021) and BaaS (Romero et al., 2021; Klimovic et al., 2018; Mahgoub et al., 2021). While this can improve control plane performance for FaaS, it sacrifices compatibility requiring cumbersome re-implementation of many cluster networking features, including authentication, DNS, and service mesh deployments for West-East traffic (Christian Posta, 2025); and limits the colocation optimizations, such as near-storage processing (Kraft et al., 2022; Amazon Web Services, 2025). The colocated deployment is fundamentally superior due to its rich feature support and vast optimization space; however, it is essential to identify and resolve the architectural bottlenecks in the existing control planes.
2.2. Overview of FaaS Control Plane Architectures
@@
To understand the bottlenecks in the current FaaS control planes, we detail the control plane architectures and operation, similar to the ones used by the leading commercial and open-source systems. Figure 2 overviews two most common serverless system architectures, both of which comprise a load balancer, a cluster manager, and worker nodes that run function instances. Load balancers and function instances constitute the data plane. Serverless function invocations arrive from the application users to the scaled-out load balancers, which route invocations to function instances for execution. If a function has no available instances to process the invocation, the cluster manager creates more instances.
@@
Control-plane architectures in state-of-the-art commercial (aws, 2025a; kna, 2025c) and open-source systems (kna, 2025a; ope, 2025a; Apache, 2025) tend to fall into two categories based on whether they create instances synchronously (on the critical path) or asynchronously (off the critical path). The synchronous approach, employed by state-of-the-art systems such as AWS Lambda (aws, 2025a) and some open-source systems (Apache, 2025), allows for an immediate scaling reaction as soon as an invocation arrives. There, the cluster manager issues an instance creation command on the critical path of an invocation to the worker node, binding the invocation for execution in the instance freshly started on that node.
@@
Other systems, including Knative (kna, 2025a), Google Cloud Run (gcr, 2025), and FunctionGraph (Liu et al., 2023), use the asynchronous approach associated with a slower reaction time: these systems typically aggregate function invocation statistics over a period of time and only start scaling when they can confirm the change in the invocation traffic. Hence, in the worst case, an invocation can wait for the entire autoscaling period (which is 2 seconds in Knative (kna, 2025b)) before the cluster manager requests an instance creation. Moreover, prior works (Liu et al., 2023; Cvetkovic et al., 2024) have found that asynchronous approaches may exhibit higher tail latencies, e.g., related to their higher queuing delays in the control plane. For example, when a single instance serves invocations, all invocations have to wait until the control plane decides and starts creating more instances.
3. Characterization of the Traffic Patterns & Existing Control Planes’ Performance
@@
In this section, we analyze serverless workload and state-of-the-art system characteristics to identify bottlenecks and optimization opportunities. We first study the traffic patterns occurring in production deployments (§3.1). Then, we study the system implications of these patterns, breaking down the scaling delays in the existing systems with synchronous and asynchronous control planes (§3.2). We also evaluate the control planes’ instance-creation throughput (§3.3) and cluster-resource usage efficiency (§3.4).
@@
We use vHive (Ustiugov et al., 2021b) research framework to configure two setups based on Knative (kna, 2025a) and Kubernetes (k8s, 2025a) for our experiments, which are widely used in commercial serverless deployments (kna, 2025c; cnc, 2020). The first configuration uses vanilla Knative that features an asynchronous control plane. The second configuration is a modified Knative version that features a synchronous control plane, with an autoscaling policy similar to AWS Lambda (Agache et al., 2020; Ustiugov et al., 2021a; aws, 2025c). §5 provides more details about these configurations and the evaluation methodology that uses sampled production traces (Microsoft Azure, 2023; Ustiugov et al., 2023).
3.1. Invocation Traffic Patterns
@@
To identify key traffic patterns and system requirements, we simulate a serverless system with a synchronous control plane with a keepalive period of 10 minutes when replaying an hour from the entire production trace (Shahrad et al., 2020; Microsoft Azure, 2023). Our simulator models the numbers of in-flight function invocations (request concurrency), active and idle instances at each moment in time, assuming instant scaling of function instances and that each instance runs on a single CPU core.
We observe that FaaS traffic is bimodal. A tiny fraction of invocations (0.1%), which we term excessive, can overload the control plane despite consuming under 2% of CPU resources. The remaining warm invocation traffic, or sustainable traffic, uses over 98% of the CPU with little impact on the control plane. This dichotomy requires a dual optimization strategy: the control plane should handle sustainable traffic with resource-efficient background scaling and placement, while optimizing for sheer scaling speed to serve latency-sensitive excessive bursts.
3.2. Analysis of the Control-Plane Scaling Delays

### v4

@@ Abstract
Serverless platforms face a trade-off: conventional cluster managers like Kubernetes offer compatibility for co-locating Function-as-a-Service (FaaS) and Backend-as-a-Service (BaaS) components of serverless applications, at the cost of high cold-start latency, whereas specialized FaaS-only systems like Dirigent achieve low latency by sacrificing compatibility, preventing integrated management and optimization. Our analysis reveals that FaaS traffic is bimodal: predictable, sustainable traffic consumes >98% of cluster resources, whereas sporadic, excessive bursts stress the control plane’s scaling latency, not its throughput.
@@
With these insights, we design PulseNet, a serverless architecture that uses a dual-track control plane tailored to both traffic types. PulseNet’s standard track manages sustainable traffic with long-lived, full-featured Regular Instances under a conventional cluster manager, preserving compatibility for the majority of the workload. To handle excessive traffic, an expedited track bypasses the slow manager to rapidly create short-lived, disposable Emergency Instances, minimizing cold-start latency and resource waste from idle instances. This hybrid approach achieves 35% better performance than Dirigent, a FaaS-only system, on a production workload at the same cost and outperforms other Kubernetes-compatible systems by 1.5–3.5, reducing the cost by up to 70%.
1 Introduction
@@
Serverless cloud applications today comprise stateless Function-as-a-Service (FaaS) and conventional stateful Backend-as-a-Service (BaaS) components, organized as DAG workflows. Developers deploy their business logic as functions while offloading the management of the underlying cloud infrastructure, for both FaaS and BaaS, entirely to the providers. The control plane is key to providing fast and timely scaling and resource efficiency when running these applications at the massive cloud scale, which can be achieved by careful placement and load balancing, and timely scaling.
@@
Unfortunately, despite the tight coupling of FaaS and BaaS components in cloud applications, their control planes lack coordination, often rendering co-design and optimization impractical. For example, AWS Lambda [agache:firecracker] and Dirigent [cvetkovic:dirigent] manage their own fleets of bare-metal nodes allocated exclusively for FaaS. Other works focus on designs for specific BaaS services [klimovic:pocket, mahgoub:sonic] used by serverless applications, such as storage and database. In contrast, a decade ago, the success of Google’s Borg cluster manager demonstrated the benefits of running various services under a single control plane, achieving better cluster utilization and reducing operational costs without compromising performance [verma:borg, tirmazi:borg]. Since then, practitioners have designed many open-source production systems [openfaas, knative, fission, nuclio] running atop Borg’s successor, Kubernetes [k8s], which features numerous optimizations and highly-optimized technologies (e.g., intelligent placement, load balancing, networking, and authentication) and can colocate any services in the same cluster. However, state-of-the-art prior works [cvetkovic:dirigent, roy:icebreaker, singhvi:atoll] have concluded that a heavyweight control plane, as in Kubernetes, is too slow to keep up with highly dynamic FaaS workloads [shahrad:serverless], calling for a full revamp of technologies already available for BaaS services running under a conventional control plane.
@@
In contrast, we show that a conventional cluster manager’s control plane, although heavyweight, can be seamlessly extended to provide sufficient scaling speed for FaaS workloads, thereby unlocking efficient co-location of FaaS and BaaS systems without compromising performance. Using production Azure Functions traces [shahrad:serverless], we present a thorough analysis of the FaaS workload and performance characteristics in systems representative of leading commercial offerings [gcr, aws-lambda-scaling] and open-source systems [openwhisk, knative_offerings]. First, we identify two kinds of invocation traffic, sustainable and excessive, showing that the former utilizes 98% of the cluster CPU resources while the latter stresses the control plane but consumes 2% of the cluster resources. Second, we demonstrate that control planes suffer from long instance-creation delays due to frequent interaction with the centralized cluster manager, e.g., when allocating cluster resources and IP addresses, and setting up network routes. However, we show that these systems’ control planes can be tuned to deliver throughput sufficient for a dynamic large-scale FaaS deployment, in contrast to the prior work [cvetkovic:dirigent] that has optimized for control-plane performance. Finally, we identify that current systems waste 70-87% of memory and 9-20% of CPU resources due to prolonged idle instance lifetime and high instance churn, respectively.
@@
Based on the obtained insights, we introduce PulseNet,111We will release the code, traces, and toolchain after publication. a serverless system designed to achieve high performance and low cost while maintaining compatibility with conventional cluster managers, such as Kubernetes [k8s]. PulseNet employs a novel dual-track control plane comprising standard and expedited tracks. The standard track is based on Knative atop Kubernetes, the industry-standard serverless cluster manager, managing long-lived, full-featured Regular Instances to handle the sustainable traffic. This track adjusts the number of instances off the critical path in cooperation with the underlying cluster manager, thereby preserving full compatibility with its rich feature set. Concurrently, the expedited track serves the excessive traffic bursts by rapidly creating Emergency Instances that completely bypass the cluster manager and its bookkeeping. This track communicates directly with a node-local agent that spawns single-use instances, which it shuts down after processing a single invocation, minimizing idle instance lifetime and associated resource waste. These disposable instances are fully compatible with FaaS workloads, albeit they forego the conventional cluster manager’s features for long-running instances, such as readiness and liveliness probes, and advanced networking capabilities, which they do not need, in exchange for much faster instance creation.
@@
We prototype PulseNet in vHive [ustiugov:benchmarking] as a seamless extension of the Knative control plane running atop unchanged Kubernetes. Evaluation with a production workload shows that PulseNet provides 35% performance improvement while maintaining cost parity with Dirigent [cvetkovic:dirigent]. Additionally, when compared with other Kubernetes-compatible alternatives, PulseNet delivers 1.5-3.5 better performance than the systems with a synchronous control plane, as in AWS Lambda [ray-serve-autoscaling], while reducing the operational cost by 8-70%; and achieves 1.7-3.5 better performance than the systems with an asynchronous control plane, as in Knative, while reducing the cost by 3-65%. Finally, compared with the systems featuring a linear regression and the state-of-the-art NHITS prediction models [challu:nhits, joosen:how], PulseNet delivers up to 4 better performance while reducing costs by 35-40%.
@@
We summarize the comparisons across state-of-the-art systems in Table 1 and illustrate the scaling speed and resource efficiency in Figure 1 on a trace snippet taken from Azure Functions trace [shahrad:serverless]. In contrast to the baselines, PulseNet closely tracks traffic changes by rapidly creating emergency instances whenever the standard track lags behind and quickly tearing them down once the traffic trend settles.
Our main contributions are:
We identify two types of invocation traffic, sustainable and excessive, with the former consuming 98% of the cluster resources and the latter stressing the control plane.
This work is the first to comprehensively characterize the performance and cost in systems with synchronous and asynchronous control planes. We also show that previously proposed intelligent scaling predictors [challu:nhits, joosen:how] incur a substantial cost overhead, with nearly 5 resource overprovisioning.
We propose PulseNet, a novel dual-track control-plane architecture that unlocks fast, resource-efficient scaling of FaaS workloads while retaining full compatibility with conventional cluster managers and their rich feature set.
PulseNet outperforms Kubernetes-compatible and FaaS-specialized systems [cvetkovic:dirigent] by 1.5-3.5 with a sampled production workload, reducing cost by up to 70%, when running synthetic and real-world benchmarks.
2 Background
Here, we describe two serverless application deployment approaches, disaggregated and co-located, and detail the state-of-the-art control-plane architectures and associated trade-offs.
2.1 FaaS&BaaS: Disaggregated or Co-Located
Serverless applications comprise functions (FaaS) along with conventional microservices (BaaS), e.g., for cross-function communication. To date, cloud providers have followed two deployment approaches: disaggregated with FaaS and BaaS components running in separate clusters with independent FaaS- [cvetkovic:dirigent, singhvi:atoll, mvondo:ofc, fuerst:faascache] and BaaS-specialized [romero:faast, klimovic:pocket, mahgoub:sonic] control planes, vs. co-located with both FaaS and BaaS orchestrated by the same cluster manager [fission, knative, openfaas, openwhisk].
Following the disaggregated approach, system architects design the cluster manager tailored for highly sporadic FaaS workloads, unlocking orders of magnitude faster scaling speed, which conventional, microservice-centric cluster managers cannot deliver [cvetkovic:dirigent]. However, despite the promised scaling-speed benefits, foregoing compatibility with the conventional cluster manager requires the cumbersome re-implementation of many essential features that academic prototypes often overlook, e.g., authentication, DNS, and service mesh deployments for West-East traffic [service_mesh], and rules out affinity-based optimizations [kraft:apiary, s3objectlambda].
In contrast, the colocation approach is fundamentally superior, as it jointly manages FaaS and BaaS components, significantly reducing overheads and cost. Two notable examples that showcase colocation benefits are Nightcore [jia:nightcore], which exploits shared memory for low-latency communication, 2 faster than with RPC, and SPRIGHT [qi:spright], which reduces the data plane’s CPU overhead with eBPF by an order of magnitude. These works rely on locality-aware scheduling, which is hard to implement across clusters. Besides, many open-source production systems adopt the colocation approach [fission, knative, openfaas, openwhisk], as it allows them to benefit from Kubernetes’ integrated management and placement optimizations [zhang:faster, bazzaz:preventing, verma:borg], support for monitoring [prometheus, grafana, opentelemetry, fluentd], network control and mounts [calico, k8s_pv], and versioned deployments [k8s_revisions].
The key obstacle to unlocking same-cluster management is designing an efficient joint control plane that meets the scaling-speed requirements of FaaS workloads while retaining the rich optimization and feature space of conventional managers like Kubernetes – which is the focus of our work.
2.2 FaaS Control-Plane Architectures
@@
To understand the bottlenecks in current FaaS control planes, we detail the control-plane architectures and operations, similar to those used by leading commercial and open-source systems. Figure 2 overviews two of the most common serverless system architectures, both of which comprise a load balancer, a cluster manager, and worker nodes that run function instances. Load balancers and function instances constitute the data plane. Serverless function invocations arrive from the application users to the scaled-out load balancers, which route invocations to function instances for execution. If a function has no available instances to process the invocation, the cluster manager creates more instances.
@@
Control-plane architectures in state-of-the-art commercial [awslambda, knative_offerings] and open-source systems [knative, openfaas, openwhisk] tend to fall into two categories based on whether they create instances synchronously (on the critical path) or asynchronously (off the critical path). The synchronous approach, employed by production systems such as AWS Lambda [awslambda] and some open-source systems [openwhisk], allows for an immediate scaling reaction as soon as an invocation arrives. There, the cluster manager issues an instance creation command on the critical path of an invocation to the worker node, binding the invocation for execution in the instance freshly started on that node.
@@
Other systems, including Knative [knative], Google Cloud Run [gcr], and FunctionGraph [liu:gap], use the asynchronous approach associated with a slower reaction time: these systems typically aggregate function invocation statistics over a period of time and only start scaling when they can confirm the change in the invocation traffic. Hence, in the worst case, an invocation can wait for the entire autoscaling period (which is 2 seconds in Knative [knative:autoscaling_period]) before the cluster manager requests an instance creation. Moreover, prior works [liu:gap, cvetkovic:dirigent] have found that asynchronous approaches may exhibit higher tail latencies, e.g., related to their higher queuing delays in the control plane. For example, when a single instance serves invocations, all invocations have to wait until the control plane decides and starts creating more instances.
3 Characterization of the Traffic Patterns & Existing Control Planes’ Performance
@@
In this section, we first study the traffic patterns occurring in production deployments (§3.1). Then, we study the system implications of these patterns, decomposing the scaling delays in existing state-of-the-art systems with synchronous and asynchronous control planes (§3.2). We also evaluate the control planes’ instance-creation throughput (§3.3) and cluster-resource usage efficiency (§3.4).
@@
We use vHive [ustiugov:benchmarking] research framework to configure two setups based on Knative [knative] and Kubernetes [k8s] for our experiments, which are widely used in commercial serverless deployments [knative_offerings, cncf:survey_2020]. The first configuration uses vanilla Knative that features an asynchronous control plane. The second configuration is a modified Knative version that features a synchronous control plane, with an autoscaling policy similar to AWS Lambda [agache:firecracker, ustiugov:stellar, aws-lambda-scaling]. Section 6 provides more details about these configurations and the evaluation methodology that uses sampled production traces [azure_trace, ustiugov:enabling].
3.1 Invocation Traffic Patterns
@@
To identify key traffic patterns and system requirements, we simulate a serverless system with a synchronous control plane, and a keep-alive period of 10 minutes when replaying an hour-long production trace [shahrad:serverless, azure_trace]. We develop a simulator that models request concurrency (i.e., the total number of in-flight function invocations) and the number of active and idle instances at each moment in time. We assume instantaneous instance scaling and that each instance runs on one CPU core.
We observe that FaaS traffic has two distinct components. The bulk of the traffic (99.9%), further referred to as sustainable, is handled by alive instances without any involvement of the control plane. Sustainable traffic uses over 98% of the CPU resources. Only the remaining 0.1% of traffic, which we term excessive, triggers the creation of new instances, causing most of the load in the control plane. Our observation corroborates the data from a major cloud provider [joosen:serverless], reporting that only 0.01% of invocations trigger cold starts.
This dichotomy motivates for a dual optimization strategy tailored to the distinct needs of each traffic component. For sustainable traffic, which dominates cluster resource consumption, the control plane must prioritize cost efficiency, leveraging features such as careful placement and background scaling to optimize cluster utilization. Conversely, excessive traffic consists of latency-sensitive bursts that require fast scaling. Crucially, excessive traffic accounts for a tiny fraction of total resources (), thus the system can trade off resource efficiency for lower latency when handling these requests without bloating overall operational costs.
3.2 Control-Plane Scaling Delays Analysis